### v1

@@ 4.5 Model deployment and validation
One Modem Two Modems Network Manual Automated Response Complaints
Disconnected Disconnected Disconnectivity Actions Taken Actions Taken Time per Month
Before LSTM 170 25 25 25 0 30 mins 20
After LSTM 20 1 1 1 170 10 sec 2
@@ 5 Discussions and future works
The deployment of the AI-driven application in the proposed cloud-native data analytics platform for an offshore O-RAN network demonstrated significant improvements in connectivity and operational efficiency. The use of LSTM models for real-time anomaly detection effectively reduced network disconnections, enhancing the user experience. These results highlight the potential of AI in managing complex network environments, particularly in challenging offshore settings.
Deploying consistent AI models across complex, multi-vendor network environments remains a significant challenge, particularly when networks span diverse regions with varying system configurations and architectures. As shown in Fig. LABEL:fig:Data1, the centralized AI models, while effective within single-vendor environments, struggle to maintain adaptability and efficiency across different network setups. This underscores the need for more flexible approaches that can ensure re-usability and privacy compliance in such diverse contexts.
To address these challenges, we propose Federated Learning (FL) as a promising solution. FL enables the training of AI models using local data from each deployment location, thus preserving data privacy and enhancing the model’s applicability across different environments. This decentralized approach ensures that AI models can be deployed consistently and efficiently, even in multi-vendor, multi-region scenarios.
Our future work will focus on implementing the FL approach across various offshore networks managed by different O-RAN vendors (Fig. LABEL:fig:Data2). By doing so, we aim to validate the effectiveness of FL in ensuring consistent performance and privacy compliance across diverse network deployments. The future work will also explore potential optimizations to further enhance the scalability and efficiency of FL in real-world applications, particularly in environments with highly heterogeneous systems and vendor setups.
@@ 6 Conclusions
RAN plays a critical role in modern telecom infrastructure, evolving towards disaggregated and open architectures like O-RAN. These innovations enable the integration of intelligent, data-driven applications to enhance network reliability and operational autonomy. However, the operation of O-RAN networks poses challenges due to immature real-world practices and complexities in managing data and applications across diverse vendor systems.
Boldyn Networks has developed a novel AI-driven cloud-native data analytics platform to address these challenges. Tested with advanced LSTM models for real-time anomaly detection, the platform significantly improves operational efficiency and enhances customer experience. Leveraging DevOps practices and tailored data lakehouse architectures for AI applications, it exemplifies sophisticated data engineering strategies.
The deployment of this platform in an offshore O-RAN network demonstrated significant improvements in connectivity and operational efficiency, validating the model’s effectiveness. However, the reliance on proprietary data highlights the need for standardized datasets to facilitate broader validation and comparison of AI models. Future research should explore the scalability of such AI-driven solutions across diverse, multi-vendor network environments. Implementing FL could ensure consistent AI model performance while preserving data privacy across different regions and system configurations.
This platform demonstrates significant potential for advancing in-RAN AI development. We aim to contribute to the community’s understanding and implementation of complex challenges in this domain, fostering innovations and improvements.
Acknowledgment
Authors Contributions

### v2

@@ 5.6 Result discussion and operational effectiveness
One Modem Two Modems Network Manual Proactive Response
Disconnected Disconnected Disconnectivity Actions Taken Actions Taken Time
Before LSTM 170 25 25 25 0 30 mins
After LSTM 20 1 1 1 170 10 sec
@@
All of this was possible because of the cloud-native data platform we built and the AI development Pyramid standards we proposed. It allowed us to bring all the data together, support the development of the AI model, and run the system smoothly. This platform also makes it easier to add more complex systems in the future when needed.
@@ 6 Discussions and future works
The deployment of the LSTM-based anomaly detection model within the Boldyn cloud-native platform has demonstrated clear benefits in enhancing connectivity, reducing service interruptions, and minimizing manual interventions. However, the current implementation is centered around a centralized training model, where raw telemetry from multiple offshore deployments is streamed to a shared processing layer for training and validation. While this model has proven effective in controlled, homogeneous environments, its scalability becomes more challenging in distributed, multi-tenant networks with differing operational policies, vendors, and data handling requirements. A key consideration for future work is the need to support diverse client ecosystems, each potentially subject to unique privacy constraints, regulatory obligations, or internal governance rules. Transferring raw data from these domains into a centralized training pipeline may not always be permissible or optimal. Additionally, deployment-specific conditions (e.g., vendor configurations, physical topologies, local anomalies) may reduce the transferability of centrally trained models, necessitating more context-aware learning mechanisms.
To address these challenges, we propose extending the platform to support the decentralized learning framework, such as federated learning (FL). As illustrated in Fig. LABEL:fig:Data1, this architecture would allow each deployment, regardless of vendor or location, to train its own local model on-site. Only model parameters or learned representations would be aggregated at the processing layer, avoiding the need to expose sensitive data. This approach is well-aligned with both data sovereignty requirements and the practical need for localized model specialization. By contrast, the current architecture (Fig. LABEL:fig:Data2) requires central collection and processing of all telemetry, which becomes less sustainable as the network footprint grows.
The federated model also offers operational efficiencies: by enabling deployments to adapt to their own evolving conditions, it reduces the frequency of centralized retraining and better accommodates real-world heterogeneity. Boldyn’s platform already includes the modular components necessary to support such a transition, namely distributed inference pipelines, secure communication layers, and model versioning capabilities.
In future work, we plan to: (1) Prototype and benchmark FL-based training across representative offshore deployments; (2) Evaluate model performance consistency across heterogeneous vendor configurations; (3) Investigate synchronization strategies and update aggregation policies under bandwidth-constrained conditions; (4) Define compliance checkpoints for deployments governed by strict data residency and audit requirements.
These steps aim to evolve the platform from a performant centralized solution to a more resilient, privacy-aligned, and scalable AI framework, which is better suited to the multi-operator, multi-region nature of emerging O-RAN ecosystems.
@@ 7 Conclusions
The radio access network (RAN) plays a critical role in modern telecom infrastructure, currently evolving towards disaggregated and open architectures such as O-RAN. These advances and innovations unlock opportunities for intelligent, data-driven applications to enhance network reliability and operational autonomy. However, the operation of O-RAN networks poses numerous research and engineering challenges due to immature real-world practices and complexities in managing data and applications across diverse vendor systems. To address these challenges, Boldyn Networks developed a novel AI-driven, cloud-native data analytics platform. Tested with advanced LSTM models for real-time anomaly detection, the platform significantly improves operational efficiency and enhances customer experience. By leveraging DevOps practices and tailored data lakehouse architectures for AI applications, the platform exemplifies sophisticated data engineering strategies. The deployment of this platform in an offshore O-RAN network demonstrated significant improvements in connectivity and operational efficiency, validating the model’s effectiveness with more than 90% of connectivity issues resolved in run-time by the LSTM model. However, the reliance on proprietary data highlights the need for standardized datasets to facilitate broader validation and comparison of AI models. Future research should explore the scalability of such AI-driven solutions across diverse, multi-vendor network environments. Implementing a decentralized platform, such as federated learning (FL) platform, could ensure consistent AI model performance while preserving data privacy across different regions and system configurations. This platform demonstrates significant potential for advancing in-RAN AI development. We aim to contribute to the community’s understanding and implementation of complex challenges in this domain, fostering innovations and improvements.
CRediT authorship contribution statement
@@
Funding
This work is a continuation of the research conducted within the Innovate UK/CELTIC-NEXT European collaborative project AIMM (AI-enabled Massive MIMO), which partially funded this research.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgment